Please answer in table form only for easy understanding. Thanks.

Diaz Company issued $148,000 face value of bonds on January 1, 2018. The bonds had a 6 percent stated rate of interest and a ten- year term. Interest is paid in cash annually, beginning December 31, 2018. The bonds were issued at 97. The straight-line method is used for amortization Required a. Use a financial statements model like the one shown below to demonstrate how (1) the January 1, 2018, bond issue and (2) the December 31, 2018, recognition of interest expense, including the amortization of the discount and the cash payment, affect the companys financial statements. Use + for increase,-for decrease, and NA for not affected. b. Determine the carrying value (face value less discount or plus premium) of the bond liability as of December 31, 2018. c. Determine the amount of interest expense reported on the 2018 income statement. d. Determine the carrying value (face value less discount or plus premium) of the bond liability as of December 31, 2019. e. Determine the amount of interest expense reported on the 2019 income statement.
Answer #1

Solution a:

Face Value of bonds = $148,000

Interest rate = 6%

Period = 10 years

Issue price = Face Value *97% = $148000*97% = $143,560

Discount on issue of bond = Face value - Issue Price = $148000 - $143560 = $4,440

Amortization of discount per year = $4440 / 10 = $444

Cash Payment for Interest = $148,000 *6% = $8,880

Interest Expense = Cash Payment for Interest + Amortization per year = $8880 + $444 = $9,324

Diaz Company
Effect of transactions on Financial Statements
Event no. Balance sheet Income Statement Statrement of Cash Flows
Assets = Liabilities + Stockholders' Equity Revenues - Expenses = Net Income
1 $1,43,560.00 $1,43,560.00 NA NA NA NA $1,43,560.00 FA
2a. NA $9,324.00 -$9,324.00 NA $9,324.00 -$9,324.00 NA NA
2b. -$8,880.00 -$8,880.00 -$8,880.00 OA

Solution b:

Carrying Value of Bond Liability on Dec 31, 2018 = Face Value - Unamortized Discount on Dec 31 2018

= $148,000 - ($4440 -$444) = $148000 - $3996 = $144,004

Solution c:

Interest Expense reported on 2018 income statement = Cash Payment for Interest + Amortization per year = $8880 +$444 = $9,324

Solution d:

Carrying Value of Bond Liability on Dec 31, 2019= Face Value - Unamortized Discount on Dec 31,2019

= $148,000 - ($3996 -$444) = $148000 - $3552 = $144,448

Solution e:

Interest Expense reported on 2019 income statement = Cash Payment for Interest + Amortization per year = $8880 +$444 = $9,324